# The Impact of Baryonic Effects on the Dynamical Masses inferred using Satellite Kinematics

### v1

@@ 2.2.4 Gas
Given the limited observational constraints on the structure of the gas in lower-mass halos, we assume that this physically motivated and self-similar description for gas in galaxy clusters remains valid across all halo mass scales. We compare our analytic gas profiles with those extracted from EAGLE halo stacks centered at $M_{\rm DMO}=10^{12.5}\>{\rm M_{\odot}}$ and $10^{13.5}\>{\rm M_{\odot}}$, finding generally good agreement (see Section 4.2). At these mass scales, adopting a larger transition radius (e.g. $x_{\rm tr}=c_{200}$ following Komatsu & Seljak 2001) yields a slightly better match to the simulated gas density profiles but has a negligible effect on the velocity dispersion at these mass scales. At higher halo masses ($M_{\rm DMO}\sim 10^{15}\>{\rm M_{\odot}}$), this alternative choice impacts the velocity dispersion with a decrease in velocity dispersion at the $0.5\%$ level (see Appendix A). We account for this effect in Section 4.3 where we explore extreme baryon models. We note that the polytropic index is largely insensitive to the choice of $x_{\rm tr}$, remaining at $\Gamma\approx 1.2$ across concentrations $4<c_{200}<7$, consistent with previous analytical and simulation results (Komatsu & Seljak 2001; Ascasibar et al. 2006; Ghirardini et al. 2017),
The enclosed mass profile for the gas component, $M_{\rm gas}(<r)$ is computed by numerically integrating this density profile. The normalization $\rho_{0}$ is set such that the total gas mass fraction within the halo corresponds to the desired value of $f_{\rm gas}$.
Figure 1: Left: Fraction of stellar mass in the diffuse stellar component as a function of halo mass, defined $f_{\mathrm{\ast,diffuse}}=M_{\rm\ast,diffuse}/(f_{\rm b}M_{\mathrm{DMO}})$, where $M_{\rm\ast,diffuse}$ is the total stellar mass within the virial radius minus the central galaxy mass. The gray scatter points represent relaxed halos in the EAGLE simulation. We fit an error-function (equation (17)) to the data. The best fit curve is shown in dashed pink. We present two extreme models that together span the shaded pink area, capturing most of the variance at a given halo mass. Middle: Fraction of ejected baryonic mass as a function of halo mass, defined as $f_{\mathrm{eject}}=M_{\rm eject}/(f_{\rm b}M_{\mathrm{DMO}})$, where $M_{\rm eject}$ is determined through equation (19). Similar to the left plot we fit this with an error-function (equation (18)), for which the best fit is shown in dashed brown and the two extreme curves cover the halo-to-halo variance of the EAGLE halos.
@@ 2.3 Adiabatic response
When gas cools and accumulates at the center of its dark matter halo, the dark matter particles will respond to the change in the gravitational potential well, causing the dark matter halo to contract. Similarly, when gas is ejected from the halo due to feedback processes, the dark matter particles becomes less bound, causing the halo to expand. In general, this response of the dark matter is difficult to model in detail. However, if the processes responsible for the change in the overall gravitational potential are sufficiently slow, the response can be assumed to be adiabatic, which is analytically tractable. In particular, if we assume all dark matter to be on circular orbits, and that the system is spherically symmetric, then $rM(<r)$, which is proportional to the square root of the specific angular momentum, is an adiabatic invariant (Blumenthal et al. 1986). Hence, under these conditions we have that
$r_{\rm f}\,M_{\rm f}(<r_{\rm f})=r_{\rm i}\,M_{\rm i}(<r_{\rm i})\,$, (22)
@@
$r_{\rm f}\,\left[M_{\rm bar}(<r_{\rm f})+M_{\rm DM,f}(<r_{\rm f})\right]=r_{\rm i}\,M_{\rm DM,i}(<r_{\rm i})\,$, (23)
where $M_{\mathrm{bar}}(<r_{\rm f})$ is the sum of the bound baryonic components (central galaxy, diffuse stars and gas). We iteratively solve this equation for $r_{\rm f}(r_{\rm i})$, from which we then compute the enclosed mass profile of the adiabatically-responded dark matter halo.
@@
where $\Upsilon=r_{\rm f}/r_{\rm i}$ is the contraction factor obtained by numerically solving equation (23), and $\nu$ is another free parameter of our model which sets the ’strength’ of the response. Setting $\nu=0$ is equivalent to no response of the dark matter, while setting $\nu=1$ preserves adiabatic invariants as described by Blumenthal et al. (1986). In general, by varying $\nu$ we can examine how strongly our results are impacted by uncertainties and complications related to how the gravitational potential responds to the galaxy formation process.
@@ 3 Model Calibration and Validation
In order to validate our baryonic model and to assure that its fiducial parameters are reasonable, we tune and compare the model against the cosmological EAGLE (’Evolution and Assembly of GaLaxies and their Environment’) simulations (Schaye et al. 2015; Crain et al. 2015). In particular, using data from both the hydrodynamical and the corresponding dark matter only runs, we examine how the various baryonic mass fractions in EAGLE scale with the DMO halo mass, and how the baryons are distributed within their host halos. We restrict our analysis to the $z=0$ simulation outputs.
@@
We follow Matthee et al. (2017) and define the stellar mass of the central galaxy as the stellar mass mass within 30 kpc, i.e., $M_{\mathrm{\ast,cen}}=M_{\ast}(<30\>{\rm kpc})$. The remaining stellar mass within $r_{200}$ (as measured in the DMO simulation) is assigned to the diffuse stellar component; $M_{\mathrm{\ast,diffuse}}=M_{\ast}(<r_{200})-M_{\mathrm{\ast,cen}}$. Similarly, the gas mass is defined as the total mass in gas inside $r_{200}$, and the ejected mass is computed using $M_{\mathrm{eject}}=f_{\rm b}\,M_{\mathrm{DMO}}-M_{\mathrm{gas}}-M_{\mathrm{\ast,cen}}-M_{\mathrm{\ast,diffuse}}$.
@@
The middle panel of Fig. 1 plots the baryonic mass fraction of the diffuse stellar component, $f_{\mathrm{\ast,diffuse}}$ as a function of $M_{\rm DMO}$. The pink-shaded region marks the 16 to 84 percentile range as inferred from EAGLE. Fitting equation (17) to this data, we obtain the relation indicated by the thick solid pink line, which has $f_{\mathrm{\ast,diffuse,max}}=0.0588$ and $\log M_{\mathrm{crit}}=12.84$ as best-fit parameters. It shows that the diffuse stellar component only comprises about one percent of the entire baryonic mass budget for halos with a DMO mass of $\sim 10^{12}\>{\rm M_{\odot}}$. This fraction steadily increases with halo mass, reaching about 6% at the mass scale of $\sim 10^{15}\mathrm{M}_{\odot}$, which exceeds the mass fraction of the central galaxy. Overall, these results are in broad agreement with observational constraints (see previous section). We therefore adopt this best-fit relation as our fiducial model for the mass fraction of the diffuse stellar component. We acknowledge that different simulations may make different predictions for $f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})$, but as we demonstrate in Appendix A, changing $f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})$ within reasonable limits has no significant impact on our results. Hence, this particular functional form does not impact the conclusions in this work.
Finally, the right-hand panel of Fig. 1 plots the ejected mass fraction of baryons, $f_{\rm eject}$, as a function of $M_{\rm DMO}$. The brown shaded region marks the 16-84 percentile range in EAGLE, while the thick solid line indicates the best-fit relation of the form of equation (18), which has $(f_{\mathrm{eject,max}},\log M_{\mathrm{char}},\alpha)=(0.661,13.17,1.79)$ as best-fit parameters. This is the $f_{\mathrm{eject}}(M_{\rm DMO})$ that we adopt for our fiducial baryonic model. Note how $f_{\rm eject}\rightarrow 0$ at the high-mass end, in agreement with the fact that the baryonic mass fractions in clusters are observed to be in agreement with the universal baryon fraction (e.g., Gonzalez et al. 2007, 2013; Dai et al. 2010; Andreon 2010; Chiu et al. 2018; Morandi et al. 2015). At $M_{\rm DMO}\sim 10^{12}\>{\rm M_{\odot}}$, though, the EAGLE simulation predicts that, on average, close to 70 percent of all baryons have been ejected. Not only does the EAGLE simulation predict a large halo-to-halo variance at the low-mass end, different cosmological simulations often make predictions for $\langle f_{\rm eject}\rangle(M_{\rm DMO})$ that are substantially different from each other due to different (subgrid) implementations of the various feedback processes at play (e.g. Ayromlou et al. 2023; Wright et al. 2024). In Section 4.3 we therefore explore the impact of alternative $\langle f_{\rm eject}\rangle(M_{\rm DMO})$ relations by varying the parameters $f_{\rm eject,max}$, $\log(M_{\mathrm{char}})$ and $\alpha$.
@@ 4 Results
Figure 2: Analytical profiles for density (left), mass (middle) for two halos of mass $M_{\rm DMO}=10^{12.5}\mathrm{M}_{\odot}$ (top) and $M_{\rm DMO}=10^{13.5}\mathrm{M}_{\odot}$ (bottom) for our fiducial model. We show the different components of the model; the total DMO profiles (black), dark matter in the baryonic model with no adiabatic response ($\nu=0$) (gray), the central galaxy (red dashed), diffuse stellar component (light pink dashed), total stellar component (red solid), gas (light blue) and total profile in the baryon model (blue). In the right figures we show the corresponding line-of-sight velocity dispersion (right) for both the DMO model (black) and the baryonic model (blue), which is obtained from their corresponding density (equiv. mass) profiles. We highlight the region where satellites are selected in Basilisk, between the minimum radius ($R_{\mathrm{min}}$ = $5\arcsec\arcsec$ due to fibre collisions) and maximum radius ($R_{\mathrm{max}}$ = 0.375 $R_{\mathrm{vir}}$) as the gray shaded area. The integrated aperture velocity dispersion is computed over the radial range $R_{\mathrm{min}},R_{\mathrm{max}}$] for both the baryonic model ($\sigma_{\mathrm{ap}}$) and DMO model ($\sigma_{\mathrm{ap,DMO}}$). The ratio of the integrated velocity dispersion between the two models is indicated in the lower left corner.
@@ 4.1 Fiducial Model
In this section, we quantify the effects of baryons on the density, mass, and velocity dispersion profiles for our fiducial model. Table 1 (right-most column) lists all adopted components. Where possible, the model is tied to empirical scaling relations, while the remaining parameters are calibrated to the EAGLE simulation. Specifically, the mass fractions of the gas and the diffuse stellar component follow the analytic fits indicated by the solid curves in Fig. 1. We also consider two cases for the halo response (see Section 2.3), namely, no response ($\nu=0$) and standard adiabatic response ($\nu=1$). For every figure in this work, we specify the value of $\nu$ used.
@@
Figure 3: Comparison between the fiducial halo model and EAGLE halos of similar mass. We show the density profiles for two halos of mass $M_{\rm DMO}=10^{12.5}\mathrm{M}_{\odot}$ (left) and $M_{\rm DMO}=10^{13.5}\mathrm{M}_{\odot}$ (right). Solid lines represent the components of our halo model: dark matter (gray), total stellar mass (central + diffuse) (red), gas (light blue), and total density profile (blue). The EAGLE results are shown as shaded regions (stacked profiles). The inset in the lower left corner provides a zoom-in on the inner dark matter density profiles, including the model with adiabatic response ($\nu=1$), shown in green.
@@
At larger radii, baryonic effects reduce the density and enclosed mass profiles. This is a result of two effects: on the one hand, some of the mass that was originally in the outskirts has condensed towards the center to build the central galaxy; on the other hand, some mass is removed from the halo entirely, as quantified by $f_{\rm eject}(M_{\rm DMO})$. The grey, vertical bands in the right-hand panels indicate $R_{\rm min}\leq R_{\rm p}\leq R_{\rm max}$, the range of projected radii over which Mitra et al. (2024) probes the line-of-sight velocities of satellite galaxies. Note that, in this range of projected radii, the differences between the line-of-sight velocity profiles of the fiducial and DMO models are small and with a negligible dependence on $R_{\rm p}$. In particular, the fact that satellites with $R_{\rm p}<R_{\rm min}=55^{\prime\prime}$ are excluded significantly reduces the sensitivity to baryonic effects, which typically manifest at smaller radii (see Section 4.4 for a more detailed discussion).

### v2

@@ 2.2.4 Gas
Given the limited observational constraints on the structure of the gas in lower-mass halos, we assume that this physically motivated and self-similar description for gas in galaxy clusters remains valid across all halo mass scales. We compare our analytic gas profiles with those extracted from EAGLE halo stacks centered at $M_{\rm DMO}=10^{12.5}\>{\rm M_{\odot}}$ and $10^{13.5}\>{\rm M_{\odot}}$, finding generally good agreement (see Section 4.2). At these mass scales, adopting a larger transition radius (e.g. $x_{\rm tr}=c_{200}$ following Komatsu and Seljak 2001) yields a slightly better match to the simulated gas density profiles but has a negligible effect on the velocity dispersion at these mass scales. At higher halo masses ($M_{\rm DMO}\sim 10^{15}\>{\rm M_{\odot}}$), this alternative choice impacts the velocity dispersion with a decrease in velocity dispersion at the $0.5\%$ level (see Appendix A). We account for this effect in Section 4.3 where we explore extreme baryon models. We note that the polytropic index is largely insensitive to the choice of $x_{\rm tr}$, remaining at $\Gamma\approx 1.2$ across concentrations $4<c_{200}<7$, consistent with previous analytical and simulation results (Komatsu and Seljak 2001; Ascasibar et al. 2006; Ghirardini et al. 2017),
The enclosed mass profile for the gas component, $M_{\rm gas}(r)$ is computed by numerically integrating this density profile. The normalization $\rho_{0}$ is set such that the total gas mass fraction within the halo corresponds to the desired value of $f_{\rm gas}$.
Figure 1: Left: Fraction of stellar mass in the central galaxy relative to the total DMO halo mass. The fiducial model follows the stellar–halo mass relation of Moster et al. (2013). For comparison, we overplot the empirical SHMRs of Moster et al. (2013), Behroozi et al. (2013) and Rodríguez-Puebla et al. (2017), as indicated, which are in excellent agreement with each other. The red shaded band marks 0.2 dex scatter around the relation of Moster et al. (2013), which, based on estimates of the scatter in the SHMR, indicates the band that is expected to enclose about 68 percent of all galaxies. The gray points indicate relaxed halos from the EAGLE simulation, which underpredict the observed relation at low halo masses. Middle: Fraction of stellar mass in the diffuse stellar component as a function of halo mass, defined as $f_{\mathrm{\ast,diffuse}}=M_{\rm\ast,diffuse}/(f_{\rm b}M_{\mathrm{DMO}})$, where $M_{\rm\ast,diffuse}$ is the total stellar mass within the virial radius minus the central galaxy mass. The solid pink curve is a fit to the EAGLE data of the form of equation (17). Right: Fraction of ejected baryonic mass as a function of halo mass, defined as $f_{\mathrm{eject}}=M_{\rm eject}/(f_{\rm b}M_{\mathrm{DMO}})$, where $M_{\rm eject}$ is determined through equation (7). The solid brown curve is a fit to the EAGLE data of the form of equation (18). The shaded bands in the middle and right panels indicate the 16th–84th percentile ranges, computed in ten equally spaced bins in halo mass.
@@ 2.3 Adiabatic response
When gas cools and accumulates at the center of its dark matter halo, the dark matter particles will respond to the change in the gravitational potential well, causing the dark matter halo to contract. Similarly, when gas is ejected from the halo due to feedback processes, the dark matter particles becomes less bound, causing the halo to expand. In general, this response of the dark matter is difficult to model in detail. However, if the processes responsible for the change in the overall gravitational potential are sufficiently slow, the response can be assumed to be adiabatic, which is analytically tractable. In particular, if we assume all dark matter to be on circular orbits, and that the system is spherically symmetric, then $rM(r)$, which is proportional to the square of the specific angular momentum, is an adiabatic invariant (Blumenthal, Faber, Flores, and Primack 1986). Hence, under these conditions we have that
$r_{\rm f}\,M_{\rm f}(r_{\rm f})=r_{\rm i}\,M_{\rm i}(r_{\rm i})\,$, (22)
@@
$r_{\rm f}\,\left[M_{\rm bar}(r_{\rm f})+M_{\rm DM,f}(r_{\rm f})\right]=r_{\rm i}\,M_{\rm DM,i}(r_{\rm i})\,$, (23)
where $M_{\mathrm{bar}}(r_{\rm f})$ is the sum of the bound baryonic components (central galaxy, diffuse stars and gas). We iteratively solve this equation for $r_{\rm f}(r_{\rm i})$, from which we then compute the enclosed mass profile of the adiabatically-responded dark matter halo.
@@
where $\Upsilon=r_{\rm f}/r_{\rm i}$ is the contraction factor obtained by numerically solving equation (23), and $\nu$ is another free parameter of our model which sets the ’strength’ of the response. Setting $\nu=0$ is equivalent to no response of the dark matter, while setting $\nu=1$ preserves adiabatic invariants as described by Blumenthal, Faber, Flores, and Primack (1986). In general, by varying $\nu$ we can examine how strongly our results are impacted by uncertainties and complications related to how the gravitational potential responds to the galaxy formation process.
@@ 3 Model Calibration and Validation
In order to validate our baryonic model and to assure that its fiducial parameters are reasonable, we tune and compare the model against the cosmological EAGLE (’Evolution and Assembly of GaLaxies and their Environment’) simulations (Schaye et al. 2015; Crain et al. 2015). In particular, using data from both the hydrodynamical and the corresponding DMO runs, we examine how the various baryonic mass fractions in EAGLE scale with the DMO halo mass, and how the baryons are distributed within their host halos. We restrict our analysis to the $z=0$ simulation outputs.
@@
We follow Matthee et al. (2017) and define the stellar mass of the central galaxy as the stellar mass mass within 30 kpc, i.e., $M_{\mathrm{\ast,cen}}=M_{\ast}(30\>{\rm kpc})$. The remaining stellar mass within $r_{200}$ (as measured in the DMO simulation) is assigned to the diffuse stellar component; $M_{\mathrm{\ast,diffuse}}=M_{\ast}(r_{200})-M_{\mathrm{\ast,cen}}$. Similarly, the gas mass is defined as the total mass in gas inside $r_{200}$, and the ejected mass is computed using equation (7).
@@
The middle panel of Fig. 1 plots the baryonic mass fraction of the diffuse stellar component, $f_{\mathrm{\ast,diffuse}}$ as a function of $M_{\rm DMO}$. The pink-shaded region marks the 16 to 84 percentile range as inferred from EAGLE. Fitting equation (17) to this data, we obtain the relation indicated by the thick solid pink line, which has $f_{\mathrm{\ast,diffuse,max}}=0.0588$ and $\log M_{\mathrm{char,diffuse}}=12.84$ as best-fit parameters. It shows that the diffuse stellar component only comprises about one percent of the entire baryonic mass budget for halos with a DMO mass of $\sim 10^{12}\>{\rm M_{\odot}}$. This fraction steadily increases with halo mass, reaching about 6% at the mass scale of $\sim 10^{15}\mathrm{M}_{\odot}$, which exceeds the mass fraction of the central galaxy. Overall, these results are in broad agreement with observational constraints (see previous section). We therefore adopt this best-fit relation as our fiducial model for the mass fraction of the diffuse stellar component. We acknowledge that different simulations may make different predictions for $f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})$, but as we demonstrate in Appendix A, changing $f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})$ within reasonable limits has no significant impact on our results. Hence, this particular functional form does not impact the conclusions in this work.
Finally, the right-hand panel of Fig. 1 plots the ejected mass fraction of baryons, $f_{\rm eject}$, as a function of $M_{\rm DMO}$. The brown shaded region marks the 16-84 percentile range in EAGLE, while the thick solid line indicates the best-fit relation of the form of equation (19), which has $(f_{\mathrm{eject,max}},\log M_{\mathrm{char,eject}},\alpha)=(0.661,13.17,1.79)$ as best-fit parameters. This is the $f_{\mathrm{eject}}(M_{\rm DMO})$ that we adopt for our fiducial baryonic model. Note how $f_{\rm eject}\rightarrow 0$ at the high-mass end, in agreement with the fact that the baryonic mass fractions in clusters are observed to be in agreement with the universal baryon fraction (e.g., Gonzalez et al. 2007, 2013; Dai et al. 2010; Andreon 2010; Chiu et al. 2018; Morandi et al. 2015). At $M_{\rm DMO}\sim 10^{12}\>{\rm M_{\odot}}$, though, the EAGLE simulation predicts that, on average, close to 70 percent of all baryons have been ejected. Not only does the EAGLE simulation predict a large halo-to-halo variance at the low-mass end, different cosmological simulations often make predictions for $\langle f_{\rm eject}\rangle(M_{\rm DMO})$ that are substantially different from each other due to different (subgrid) implementations of the various feedback processes at play (e.g. Ayromlou et al. 2023; Wright et al. 2024). In Section 4.3 we therefore explore the impact of alternative $\langle f_{\rm eject}\rangle(M_{\rm DMO})$ relations by varying the parameters $f_{\rm eject,max}$, $\log(M_{\mathrm{char,eject}})$ and $\alpha$.
@@ 4 Results
Figure 2: Left: Analytical density profiles for two halos of mass $M_{\rm DMO}=10^{12.5}\mathrm{M}_{\odot}$ (top) and $M_{\rm DMO}=10^{13.5}\mathrm{M}_{\odot}$ (bottom) for our fiducial model. We show the different components of the model; the total DMO profiles (black), dark matter in the baryonic model with no adiabatic response ($\nu=0$) (gray), the central galaxy (red dashed), diffuse stellar component (light pink dashed), total stellar component (red solid), gas (light blue) and total profile in the baryon model (blue). Middle: The corresponding enclosed mass profiles for the same two halos. Right: The line-of-sight velocity dispersion for both the DMO model (black) and the baryonic model (blue), which are obtained from their corresponding density (equiv. mass) profiles. We highlight the region where satellites are selected in Basilisk, between the minimum radius ($R_{\mathrm{min}}$ = $5\arcsec\arcsec$ due to fibre collisions) and maximum radius ($R_{\mathrm{max}}$ = 0.375 $R_{\mathrm{vir}}$) as the gray shaded area. The integrated aperture velocity dispersion is computed over the radial range $R_{\mathrm{min}},R_{\mathrm{max}}$] for both the baryonic model ($\sigma_{\mathrm{ap}}$) and DMO model ($\sigma_{\mathrm{ap,DMO}}$). The ratio of the integrated velocity dispersion between the two models is indicated in the lower left corner.
@@ 4.1 Fiducial Model
In this section, we quantify the effects of baryons on the density, mass, and velocity dispersion profiles for our fiducial model. Table 1 (right‑most column) lists all adopted components. Where possible, the model is tied to empirical scaling relations, while the remaining parameters are calibrated to the EAGLE simulation. Specifically, the mass fractions of the ejected baryons and the diffuse stellar component follow the analytic fits indicated by the solid curves in Fig. 1. We also consider two cases for the halo response (see Section 2.3), namely, no response ($\nu=0$) and standard adiabatic response ($\nu=1$). For every figure in this work, we specify the value of $\nu$ used.
@@
Figure 3: Comparison between the fiducial halo model and EAGLE halos of similar mass. We show the density profiles for two halos of mass $M_{\rm DMO}=10^{12.5}\mathrm{M}_{\odot}$ (left) and $M_{\rm DMO}=10^{13.5}\mathrm{M}_{\odot}$ (right). Solid lines represent the components of our halo model: dark matter (gray), total stellar mass (central + diffuse) (red), gas (light blue), and total density profile (blue). To quantify the halo-to-halo variation in EAGLE halos, we compute the central 95% spread in density in bins of $r/r_{200}$ for two stacks of EAGLE halos whose median DMO masses match the model: $10^{12.25}<M_{\rm DMO}/\>{\rm M_{\odot}}<10^{12.85}$ (480 halos, left panel) and $10^{13.3}<M_{\rm DMO}/\>{\rm M_{\odot}}<10^{13.9}$ (31 halos, right panel). This spread is shown as shaded regions in matching colors for each component. In the main panels, the analytical model does not include halo response ($\nu=0$). The insets in the lower-left corner zoom in on the inner dark matter density profiles: the gray curve shows the $\nu=0$ model (as in the main panel), while the green curve includes adiabatic response ($\nu=1$).
@@
At larger radii, baryonic effects slightly reduce the density and enclosed mass profiles. This is a result of two effects: on the one hand, some of the mass that was originally in the outskirts has condensed towards the center to build the central galaxy; on the other hand, some mass is removed from the halo entirely, as quantified by $f_{\rm eject}(M_{\rm DMO})$. The grey, vertical bands in the right-hand panels indicate $R_{\rm min}\leq R_{\rm p}\leq R_{\rm max}$, the range of projected radii over which Mitra et al. (2024) probes the line-of-sight velocities of satellite galaxies. Note that, in this range of projected radii, the differences between the line-of-sight velocity profiles of the fiducial and DMO models are small and with a negligible dependence on $R_{\rm p}$. In particular, the fact that satellites with $R_{\rm p}<R_{\rm min}=55^{\prime\prime}$ are excluded significantly reduces the sensitivity to baryonic effects, which typically manifest at smaller radii (see Section 4.4 for a more detailed discussion).